# Learning Dense Flow Field for Highly-accurate Cross-view Camera Localization

### v1

@@ Abstract
This paper addresses the problem of estimating the 3-DoF camera pose for a ground-level image with respect to a satellite image that encompasses the local surroundings. We propose a novel end-to-end approach that leverages the learning of dense pixel-wise flow fields in pairs of ground and satellite images to calculate the camera pose. Our approach differs from existing methods by constructing the feature metric at the pixel level, enabling full-image supervision for learning distinctive geometric configurations and visual appearances across views. Specifically, our method employs two distinct convolution networks for ground and satellite feature extraction. Then, we project the ground feature map to the bird’s eye view (BEV) using a fixed camera height assumption to achieve preliminary geometric alignment. To further establish content association between the BEV and satellite features, we introduce a residual convolution block to refine the projected BEV feature. Optical flow estimation is performed on the refined BEV feature map and the satellite feature map using flow decoder networks based on RAFT. After obtaining dense flow correspondences, we apply the least square method to filter matching inliers and regress the ground camera pose. Extensive experiments demonstrate significant improvements compared to state-of-the-art methods. Notably, our approach reduces the median localization error by 89%, 19%, 80% and 35% on the KITTI, Ford multi-AV, VIGOR and Oxford RobotCar datasets, respectively.
@@ 1 Introduction
Our key idea is to learn dense flow fields between cross-views and directly compute accurate ground camera poses from such flows in an end-to-end manner. To achieve this, we develop two distinct convolutional neural networks for feature extraction from the aerial and ground images, respectively. But ground and aerial feature maps exhibit significant domain differences in terms of spatial configuration and visual appearance, posing a critical challenge for the neural network to learn effective matching on such feature maps. Furthermore, due to the absence of depth information in the ground images, direct cross-view matching leads to a one-to-many correspondence issue. For instance, a building facade in the ground image may correspond to multiple edges in the aerial image. Existing methods often assume a fixed ground height to enable the projection from the ground to the satellite imageshi2022beyond . In line with this approach, our method employs the same strategy to geometrically project the ground features, obtaining a bird’s eye view (BEV) feature map that corresponds to the ground image. This projection facilitates spatial alignment between the cross-view feature maps. As mentioned above, for the uncertainty in the ground height assumption, we introduce a RefineBlock to further refine the BEV features and achieve content alignment. Subsequently, optical flow estimation is performed on the refined BEV feature map and the satellite feature map using flow decoder networks based on RAFTteed2020raft . Specifically, we calculate multi-scale correlation volumes using the BEV and satellite features, and employ GRU modules iteratively for dense flow estimation. Furthermore, we assign scores to each established matching relationship based on their similarity, effectively reducing the weight of erroneous matches. A notable advantage of computing flows on the BEV and satellite features is that the matching relationship allows us to directly utilize the least square method to solve the 3 DoF pose, i.e., the 2 DoF translation, and 1 DoF rotation.In order to learn robust matching results, we use the dense optical flow as well as the ground truth pose as supervision to train the network progressively.
@@ 2 Related Works
Image retrieval methods are widely utilized for same-view localization tasksradenovic2018fine ; sarlin2020superglue ; arandjelovic2016netvlad ; perronnin2010large . This approach assumes the availability of a reference dataset containing a vast number of geotagged ground images. Query images are then matched with ground images in the reference dataset to determine their corresponding geographic locations. However, this method requires a significant number of ground markers, which can be expensive. In cross-view retrieval tasks, a common approach is to create separate global descriptors for the ground image and the satellite image, and then compare their similarityhu2018cvm ; shi2019spatial ; yang2021cross ; toker2021coming ; vo2016localizing .Alternatively, recent studies have explored fusing consecutive multi-frame ground images to generate a descriptor, which is then matched with the descriptor of the satellite imagevyas2022gama ; shi2023cvlnet . In these cross-view retrieval methods, the camera is assumed to be located at the center of the satellite map. However, this assumption limits the generalization of these methods.
@@
This work differs from the retrieval task in that it relies on an initial estimate of a given camera pose and then proceeds to refine it to obtain a pose that is close to the true value. Previous works have used various approaches to tackle this problem. For example, VIGORzhu2021vigor first uses an end-to-end framework to roughly locate the query through retrieval, and then refines the location by predicting offsets through regression. Xia et al.xia2022visual formulates the localization problem as a multiclass classification problem and uses a dense multimodal space for localization. shi2022beyond uses a geometric projection module to align the cross-view features, and then optimize the camera pose end-to-end using the Levenberg-Marquardt algorithmlevenberg1944method . Wang et al.wang2022satellite uses 3D points to bridge the geometric gap between ground and overhead views, iteratively fusing ground images with 3D points iteratively for pose estimation. SliceMatchlentsch2022slicematch is given a set of candidate photo poses and then compares a single ground descriptor with a set of pose dependent satellite map descriptors to generate the final camera pose. It is worth noting that previous works have primarily relied on global descriptors zhu2021vigor ; xia2022visual or local descriptorsshi2022beyond ; levenberg1944method ; wang2022satellite ; lentsch2022slicematch for cross-view positional estimation. However, in our work, we propose an alternative approach that relies on dense point matching, which we believe to be a reliable basis for cross-view localization.
@@ 3.2 Dense Flow Estimation
Our architecture consists of two separate feature extractors, namely the feature encoder and the context encoder. Here, we have utilized a feature encoder equipped with view unification module to deal with the significant differences in viewing angles between the $\textbf{I}_{g}$ and $\textbf{I}_{s}$. The context encoder has the same architecture as the feature encoder but is designed to extract only semantic and contextual information from the reference image $\textbf{I}_{g}$.
@@ 4.1 Datasets
We conducted tests on the KITTIgeiger2013vision ; shi2022beyond , Ford multi-AVAgarwal_Vora_Pandey_Williams_Kourous_McBride_2020 ; shi2022beyond , VIGORzhu2021vigor ; lentsch2022slicematch , and Oxford RobotCarmaddern20171 ; maddern2020real ; xia2022visual datasets. The query images in KITTI, Ford multi-AV, and Oxford RobotCar datasets have a limited horizontal field of view (HFoV), while the query images in the VIGOR dataset are panoramic. Given that KITTI and VIGOR were captured on different road segments, we evaluated the model’s performance on both same area and cross area scenarios using these two datasets. Ford multi-AV was mainly captured in suburban areas with fewer reference points, and Oxford RobotCar was captured on the same road segment at different times. We tested the model’s generalization ability to different environments, including suburban scenes, varying lighting conditions, and road situations, on these two datasets. The datasets used in this paper are obtained under academic licenses and are not original datasets specifically created for this work. A more detailed description of each dataset will be provided in the appendix.
@@ 4.2 Evaluation Metrics
We followed the settings of lentsch2022slicematch ; shi2022beyond ; xia2022visual . Specifically, we calculate the mean and median errors in meters between predicted and ground truth positions, as well as the mean and median angular errors in degrees between predicted and ground truth bearings. For the KITTI and Ford multi-AV dataset, we also included the recall values for longitudinal and lateral localization errors, as well as orientation estimation errors, under certain thresholds. Our localization thresholds were set at 1 meter and 5 meters, while our orientation estimation thresholds were set at 1 degree and 5 degrees.
@@ 4.3 Implementation Details
4.4 Comparison with other methods
Same-Area
We conducted performance evaluations of the model in the same region on both the KITTI and Vigor datasets. The results of the KITTI dataset tests are shown in Table 1. Our method significantly outperformed previous algorithms in camera pose estimation, with an average error reduction of 81% and a median error reduction of 89% compared to SliceMatch. Notably, our algorithm’s predictive performance for vertical position in the same region was significantly higher than that of other algorithms, indicating that despite the limited field of view, our approach learned the semantic relationship between roadside buildings in this area and their corresponding features in satellite images. In Table 3, the testing results on the VIGOR dataset show that our method achieves significantly lower average and median localization errors than SliceMatch by 42% and 62%, respectively, when the direction is known. When the direction is unknown, our algorithm still performs well, with both the average and median localization errors lower than those of SliceMatch. Furthermore, our algorithm demonstrates good performance in angle prediction, with average and median angle errors lower than those of SliceMatch by 67% and 69%, respectively.
@@ 4.4 Comparison with other methods
Cross-Area
Generalizing new ground images across different regions is a more challenging task than within the same region, as the test area may appear vastly different from the training area. Despite this increased difficulty, our approach still demonstrates strong performance. As shown in Table 1 Cross-Area, Our average localization error is 40% lower than SliceMatch, and our median localization error is 64% lower than SliceMatch. Compared to descriptor-based matching methods, our approach exhibits greater robustness. In the cross area of VIGOR, our algorithm performs better than previous methods in both known and unknown directions. Additionally, as shown in Figure 3, we tested the performance of our algorithm across various angles using a model trained on unknown directions. Our algorithm demonstrated strong stability against noise for different angle variations in both same area and cross area.
Vary-Environment
In addition, to validate the performance of our algorithm under different environmental conditions, we tested it on the Ford multi-AV dataset and the Oxford RobotCar dataset. The results of the Ford multi-AV dataset are presented in Table 2, which is primarily focused on suburban scenes with fewer surrounding buildings and difficulty in obtaining semantic information, resulting in generally poor localization accuracy. Moreover, in the limited HFOV scenario, the longitudinal position is more challenging than the lateral position in the driving direction. Consequently, the recall rate for the longitudinal position is significantly lower than that of the lateral position, and this trend applies to all the compared methods. Nonetheless, our algorithm still outperforms previous algorithms on this dataset, and we found that our lateral recall rate is significantly higher, indicating our algorithm’s ability to capture road surface information under semantically unclear conditions. And our algorithm demonstrated superior performance across the three distinct test sets of the Oxford RobotCar dataset, as shown in Table 4. These results suggest that our approach exhibits strong generalization capabilities across diverse environmental conditions, seasonal variations, and roadway scenarios.
@@ 5 Conclusion
This paper presents a novel approach based on dense matching to address the problem of cross-view localization. Our approach utilizes view unification module to bridge the visual gap between satellite and ground images. Additionally, we introduce a matching loss to provide stronger supervision for the model during training. To evaluate the effectiveness of our algorithm, we conduct experiments on four datasets that vary in geographical regions, environmental conditions, and road scenarios. Our results demonstrate impressive generalization capability of our approach. In future work, we plan to build upon this foundation by exploring the integration of spatial and temporal information to further improve the accuracy and robustness of our approach for cross-view localization.

### v2

@@ Abstract
This paper addresses the problem of estimating the 3-DoF camera pose for a ground-level image with respect to a satellite image that encompasses the local surroundings. We propose a novel end-to-end approach that leverages the learning of dense pixel-wise flow fields in pairs of ground and satellite images to calculate the camera pose. Our approach differs from existing methods by constructing the feature metric at the pixel level, enabling full-image supervision for learning distinctive geometric configurations and visual appearances across views. Specifically, our method employs two distinct convolution networks for ground and satellite feature extraction. Then, we project the ground feature map to the bird’s eye view (BEV) using a fixed camera height assumption to achieve preliminary geometric alignment. To further establish the content association between the BEV and satellite features, we introduce a residual convolution block to refine the projected BEV feature. Optical flow estimation is performed on the refined BEV feature map and the satellite feature map using flow decoder networks based on RAFT. After obtaining dense flow correspondences, we apply the least square method to filter matching inliers and regress the ground camera pose. Extensive experiments demonstrate significant improvements compared to state-of-the-art methods. Notably, our approach reduces the median localization error by 89%, 19%, 80%, and 35% on the KITTI, Ford multi-AV, VIGOR, and Oxford RobotCar datasets, respectively.
@@ 1 Introduction
Our key idea is to learn dense flow fields between cross-views and directly compute accurate ground camera poses from such flows in an end-to-end manner. To achieve this, we develop two distinct convolutional neural networks for feature extraction from the aerial and ground images, respectively. But ground and aerial feature maps exhibit significant domain differences in terms of spatial configuration and visual appearance, posing a critical challenge for the neural network to learn effective matching on such feature maps. Furthermore, due to the absence of depth information in the ground images, direct cross-view matching leads to a one-to-many correspondence issue. For instance, a building facade in the ground image may correspond to multiple edges in the aerial image. Existing methods often assume a fixed ground height to enable the projection from the ground to the satellite imageshi2022beyond . In line with this approach, our method employs the same strategy to geometrically project the ground features, obtaining a bird’s eye view (BEV) feature map that corresponds to the ground image. This projection facilitates spatial alignment between the cross-view feature maps. As mentioned above, for the uncertainty in the ground height assumption, we introduce a RefineBlock to further refine the BEV features and achieve content alignment. Subsequently, optical flow estimation is performed on the refined BEV feature map and the satellite feature map using flow decoder networks based on RAFTteed2020raft . Specifically, we calculate multi-scale correlation volumes using the BEV and satellite features and employ GRU modules iteratively for dense flow estimation. Furthermore, we assign scores to each established matching relationship based on their similarity, effectively reducing the weight of erroneous matches. A notable advantage of computing flows on the BEV and satellite features is that the matching relationship allows us to directly utilize the least square method to solve the 3 DoF pose, i.e., the 2-DoF translation, and 1-DoF rotation. In order to learn robust matching results, we use the dense optical flow as well as the ground truth pose as supervision to train the network progressively.
@@ 2 Related Works
Image retrieval methods are widely utilized for same-view localization tasksradenovic2018fine ; sarlin2020superglue ; arandjelovic2016netvlad ; perronnin2010large . This approach assumes the availability of a reference dataset containing a vast number of geotagged ground images. Query images are then matched with ground images in the reference dataset to determine their corresponding geographic locations. However, this method requires a significant number of ground markers, which can be expensive. As satellite maps become more widely available, cross-view localization has gained attention as a potential alternative. In cross-view retrieval tasks, a common approach is to create separate global descriptors for the ground image and the satellite image, and then compare their similarityhu2018cvm ; shi2019spatial ; yang2021cross ; toker2021coming ; vo2016localizing . Alternatively, recent studies have explored fusing consecutive multi-frame ground images to generate a descriptor, which is then matched with the descriptor of the satellite imagevyas2022gama ; shi2023cvlnet . In this cross-view retrieval methods, the camera is assumed to be located at the center of the satellite map. However, this assumption limits the generalization of these methods.
@@
This work differs from the retrieval task in that it relies on an initial estimate of a given camera pose and then proceeds to refine it to obtain a pose that is close to the true value. Previous works have used various approaches to tackle this problem. For example, VIGORzhu2021vigor first uses an end-to-end framework to roughly locate the query through retrieval and then refines the location by predicting offsets through regression. Xia et al.xia2022visual formulates the localization problem as a multi-class classification problem and uses a dense multimodal space for localization. shi2022beyond uses a geometric projection module to align the cross-view features, and then optimize the camera pose end-to-end using the Levenberg-Marquardt algorithmlevenberg1944method . Wang et al.wang2022satellite uses 3D points to bridge the geometric gap between ground and overhead views, iteratively fusing ground images with 3D points iteratively for pose estimation. SliceMatchlentsch2022slicematch is given a set of candidate photo poses and then compares a single ground descriptor with a set of pose-dependent satellite map descriptors to generate the final camera pose. It is worth noting that previous works have primarily relied on global descriptors zhu2021vigor ; xia2022visual or local descriptorsshi2022beyond ; levenberg1944method ; wang2022satellite ; lentsch2022slicematch for cross-view positional estimation. However, in our work, we propose an alternative approach that relies on dense point matching, which we believe to be a reliable basis for cross-view localization.
@@ 3.2 Dense Flow Estimation
Our architecture consists of two separate feature extractors, namely the feature encoder and the context encoder. Here, we have utilized a feature encoder equipped with a view unification module to deal with the significant differences in viewing angles between the $\textbf{I}_{g}$ and $\textbf{I}_{s}$. The context encoder has the same architecture as the feature encoder but is designed to extract only semantic and contextual information from the reference image $\textbf{I}_{g}$.
@@ 4.1 Datasets
We conducted tests on the KITTIgeiger2013vision ; shi2022beyond , Ford multi-AVAgarwal_Vora_Pandey_Williams_Kourous_McBride_2020 ; shi2022beyond , VIGORzhu2021vigor ; lentsch2022slicematch , and Oxford RobotCarmaddern20171 ; maddern2020real ; xia2022visual datasets. The query images in KITTI, Ford multi-AV, and Oxford RobotCar datasets have a limited horizontal field of view (HFoV), while the query images in the VIGOR dataset are panoramic. Given that KITTI and VIGOR were captured on different road segments, we evaluated the model’s performance on both same-area and cross-area scenarios using these two datasets. Ford multi-AV was mainly captured in suburban areas with fewer reference points, and Oxford RobotCar was captured on the same road segment at different times. We tested the model’s generalization ability to different environments, including suburban scenes, varying lighting conditions, and road situations, on these two datasets. The datasets used in this paper are obtained under academic licenses and are not original datasets specifically created for this work. A more detailed description of each dataset will be provided in the appendix.
@@ 4.2 Evaluation Metrics
We followed the settings of lentsch2022slicematch ; shi2022beyond ; xia2022visual . Specifically, we calculate the mean and median errors in meters between predicted and ground truth positions, as well as the mean and median angular errors in degrees between predicted and ground truth bearings. For the KITTI and Ford multi-AV datasets, we also included the recall values for longitudinal and lateral localization errors, as well as orientation estimation errors, under certain thresholds. Our localization thresholds were set at 1 meter and 5 meters, while our orientation estimation thresholds were set at 1 degree and 5 degrees.
@@ 4.3 Implementation Details
4.4 Same-Area Evaluations
We conducted performance evaluations of the model in the same region on both the KITTI and Vigor datasets. The results of the KITTI dataset tests are shown in Table 1. Our method significantly outperformed previous algorithms in camera pose estimation, with an average error reduction of 81% and a median error reduction of 89% compared to SliceMatch. Notably, our algorithm’s predictive performance for vertical position in the same region was significantly higher than that of other algorithms, indicating that despite the limited field of view, our approach learned the semantic relationship between roadside buildings in this area and their corresponding features in satellite images.
In Table 3, the testing results on the VIGOR dataset show that our method achieves significantly lower average and median localization errors than SliceMatch by 42% and 62%, respectively, when the direction is known. When the direction is unknown, our algorithm still performs well, with both the average and median localization errors lower than those of SliceMatch.
Furthermore, our algorithm demonstrates good performance in angle prediction, with average and median angle errors lower than those of SliceMatch by 67% and 69%, respectively.
@@ 4.4 Same-Area Evaluations
4.5 Cross-Area Generalization
Generalizing new ground images across different regions is a more challenging task than within the same region, as the test area may appear vastly different from the training area. Despite this increased difficulty, our approach still demonstrates strong performance. As shown in Table 1 Cross-Area, Our average localization error is 40% lower than SliceMatch, and our median localization error is 64% lower than SliceMatch. Compared to descriptor-based matching methods, our approach exhibits greater robustness.
In the cross area of VIGOR, our algorithm performs better than previous methods in both known and unknown directions. Additionally, as shown in Figure 3, we tested the performance of our algorithm across various angles using a model trained in unknown directions. Our algorithm demonstrated strong stability against noise for different angle variations in both the same area and cross area.
4.6 Vary-Environment Generalization
In addition, to validate the performance of our algorithm under different environmental conditions, we tested it on the Ford multi-AV dataset and the Oxford RobotCar dataset. The results of the Ford multi-AV dataset are presented in Table 2, which is primarily focused on suburban scenes with fewer surrounding buildings and difficulty in obtaining semantic information, resulting in generally poor localization accuracy. Moreover, in the limited HFOV scenario, the longitudinal position is more challenging than the lateral position in the driving direction. Consequently, the recall rate for the longitudinal position is significantly lower than that of the lateral position, and this trend applies to all the compared methods. Nonetheless, our algorithm still outperforms previous algorithms on this dataset, and we found that our lateral recall rate is significantly higher, indicating our algorithm’s ability to capture road surface information under semantically unclear conditions. Our algorithm demonstrated superior performance across the three distinct test sets of the Oxford RobotCar dataset, as shown in Table 4. These results suggest that our approach exhibits strong generalization capabilities across diverse environmental conditions, seasonal variations, and roadway scenarios.
@@ 4.7 Ablation Study
In Figure 5, we conducted a qualitative analysis of the confidence mechanism on the KITTI dataset. We utilized a threshold of 0.5 to distinguish between high-confidence matches and low-confidence matches. In these visualizations, circles represent matching points with high confidence, while stars indicate matching points with low confidence. Additionally, for clarity, we displayed estimated corresponding points on both the original ground and satellite images. From the visualizations, it is evident that pixels in the bird’s-eye-view image, which are mutually visible in both ground and satellite images (e.g., road areas), tend to have higher confidence. On the other hand, pixels that are visible only in one view (such as building rooftops, tree canopies, and occluded regions) generally exhibit lower confidence. As a result, our confidence map effectively selects reliable matching points for accurate pose estimation. This holds significant implications for improving precision.
@@ 5 Conclusion
This paper presents a novel approach based on dense matching to address the problem of cross-view localization. Our approach utilizes a view unification module to bridge the visual gap between satellite and ground images. Additionally, we introduce a matching loss to provide stronger supervision for the model during training. To evaluate the effectiveness of our algorithm, we conduct experiments on four datasets that vary in geographical regions, environmental conditions, and road scenarios. Our results demonstrate the impressive generalization capability of our approach. One possible future direction is to explore the integration of spatial and temporal information, to further improve the accuracy and robustness of our approach for cross-view localization.
Refer to caption Ground image
Refer to caption Sat. image
Refer to caption Confidence map
Refer to caption BEV feature map Sat. feature map
Refer to caption Ground image
Refer to caption Sat. image
Refer to caption Confidence map
Refer to caption BEV feature map Sat. feature map
Figure 5: Matching confidence illustration. The top row of each subfigure shows the ground and satellite images, highlighting corresponding points. Circles indicate confident correspondences (score > 0.5), while stars represent potentially incorrect correspondences (score < 0.5). Matching points share the same color and shape in both images. In the bottom row, the left image shows confidence scores on the BEV feature map (blue for > 0.5, green for < 0.5). The middle and right images depict the BEV and satellite feature maps, respectively, with connected lines indicating matching pixels. The circles and stars in these images align with the corresponding points in the top row.
6 Acknowledgement
This work was supported in part by the National Natural Science Foundation of China (No. 62302220), in part by the China Postdoctoral Foundation (No. 2023M731691), and in part by the Jiangsu Funding Program for Excellent Postdoctoral Talent (No.2022ZB268).